How much water must be added to 300l of a solution so that it's concentration may change from 25% to 20%
Answer by stanbon(75887)

How much water must be added to 300l of a solution so that it's concentration may change from 25% to 20%
------
Equation::
water + water = water
0.75*3001 + 1*x = 0.80(3001+x)
--------
75*3001 + 100x = 80*3001 + 80x
-------
20x = 5*3001
x = 750.25 units of water must be added
=============
Cheers,
Stan H.
-------------